OpenStudy (anonymous):

A jar containing 2011 fair coins also contains one tricked coin (it has heads on both sides). A coin is randomly selected from this jar and flipped 2012 times. It comes up heads each time. What is the probability that this was one of the fair coins?

OpenStudy (rsadhvika):

\((2010/2011)*(1/2^{2012})\) ?

OpenStudy (anonymous):

It's supposed to be that, but how do you get that?

OpenStudy (rsadhvika):

(Probability of picking fair coin, while selecting randomly) * (probability of head) * 2012

OpenStudy (rsadhvika):

(Probability of picking fair coin, while selecting randomly) * (probability of head) ^ 2012

OpenStudy (anonymous):

Oh ok. So: Probability of picking fair coin = \(\frac{2011}{2012}\) Probability of heads = \(\frac 12\) That's how it works?

OpenStudy (rsadhvika):

yes

OpenStudy (anonymous):

Alright. Why is it to the 2012 power though? I get that's the max number of coins, but why would it be raised to it?

OpenStudy (rsadhvika):

you are doing the experiment 2012 times P(first time heads) = 1/2 P(first two times heads) = 1/2 * 1/2 P(first n times heads) = 1/2 * 1/2 *...... n times

OpenStudy (rsadhvika):

tossing a coin has two possible outcomes each has a probability of 1/2

OpenStudy (rsadhvika):

if you perform the experiment n times probability is 1/2^n

OpenStudy (anonymous):

Oh. I see. Thank you :)

OpenStudy (rsadhvika):

welcome :)